One molecule of a substance absorbs one quantum of energy. Find the enerhy involved wheb 1.5 mole of the sunbstance absorvs red light of frequency 7.5 * 1014.

Grade:

1 Answers

Harsh
4 years ago
given :- frequency=7.5 *10^14 then
one molecule absorbe 1.5 mole of photon
 
so,
     Et = n * Ep
     Et = 1.5*6.63*3*10^-34 / 7.5*10^14
 
      Et = 4.49 * 10^5